The residents of Vegopia spend all of their income on cauliflower, broccoli, and carrots. In 2016, they spend a total of $210 for 100 heads of cauliflower, $70 for 50 bunches of broccoli, and $45 for 450 carrots. In 2017, they spend a total of $225 for 90 heads of cauliflower, $112 for 80 bunches of broccoli, and $120 for 400 carrots.
Complete the following table by calculating the price of one unit of each vegetable in each year.
Year
Cauliflower
Broccoli
Carrots
2016
2017
Using 2016 as the base year, the CPI for 2016 is , and the CPI for 2017 is .
The inflation rate in 2017 isusing the CPI.
1
Expert's answer
2018-09-18T09:59:08-0400
In 2016, they spend a total of $210 for 100 heads of cauliflower, so Pcaul = $2.1, $70 for 50 bunches of broccoli, so Pb = $1.4 and $45 for 450 carrots, so Pcar = $0.1.
In 2017, they spend a total of $225 for 90 heads of cauliflower, so Pcaul = $2.5, $112 for 80 bunches of broccoli, so Pb = $1.4, and $120 for 400 carrots, so Pcar = $0.3.
Year Cauliflower Broccoli Carrots
2016 2.1 1.4 0.1
2017 2.5 1.4 0.3
Using 2016 as the base year, the CPI for 2016 is 100, and the CPI for 2017 is (225 + 112 + 120)/(210 + 70 + 45)*100 = 141.
The inflation rate in 2017 is using the CPI is:
Inflation rate = (141 - 100)/100*100% = 41%.